### v3

# MAP: Low-compute Model Merging with Amortized Pareto Fronts via Quadratic Approximation

Abstract
Model merging has emerged as an effective approach to combine multiple single-task models, fine-tuned from the same pre-trained model, into a multitask model. This process typically involves computing a weighted average of the model parameters without any additional training. Existing model-merging methods focus on enhancing average task accuracy. However, interference and conflicts between the objectives of different tasks can lead to trade-offs during model merging. In real-world applications, a set of solutions with various trade-offs can be more informative, helping practitioners make decisions based on diverse preferences. In this paper, we introduce a novel and low-compute algorithm, Model Merging with Amortized Pareto Front (MAP). MAP efficiently identifies a Pareto set of scaling coefficients for merging multiple models to reflect the trade-offs. MAP approximates the evaluation metrics of the various tasks using a quadratic approximation surrogate model derived from a pre-selected set of scaling coefficients, enabling amortized inference. Experimental results on vision and natural language processing tasks show that MAP can accurately identify the Pareto front. To further reduce the required computation of MAP, we propose (1) a Bayesian adaptive sampling algorithm and (2) a nested merging scheme with multiple stages.
1 Introduction
Large pre-trained foundation models have become available in many real-world applications [60; 54; 13]. This increasing availability has led to a popular practice of fine-tuning these pre-trained models to adapt to a wide range of downstream tasks. Practitioners can independently fine-tune the same pre-trained model, such as CLIP style models [45; 63; 69], large language models [6; 47; 55; 28], etc., and then release the fine-tuned models without releasing the training data. As the deployment of such fine-tuned models increases, combining models with identical architectures and initializations has emerged as a promising approach to combine their respective capabilities. This is useful, especially in scenarios where the training data for each task is private and cannot be shared, such as individual-level patient data in a hospital and behavior data in social media recommendation systems.
Existing methods for merging models typically involve calculating a weighted average of the parameters from multiple models to enhance performance uniformly across various tasks. However, this approach often overlooks the conflicts among the diverse objectives of these tasks, which can lead to trade-offs in terms of model performance on various tasks. In real-world applications, it is often useful to obtain a set of Pareto optimal solutions rather than a single model. Such solutions allow practitioners to choose among different trade-offs, depending on their specific needs. For example, hospitals specializing in certain areas might prefer a model that excels in tasks relevant to their specialty while maintaining adequate performance across a broader spectrum of diseases.
[…]
Figure 1: The overall process of MAP. Step 1: select 2 tasks and their corresponding task vectors. Step 2: sample a few scaling weights $\mathbf{c}$ to query the task metrics accordingly. Step 3: use the Quadratic model to approximate the $\mathbf{c}\rightarrow\textit{metrics}$ mapping as a surrogate model. Step 4: Use NSGA-III to find amortized Pareto fronts. Figure (a) shows a contour plot of the actual accuracy landscape for the ViT models [17] obtained from 100 scaling coefficients (sampled uniformly) evaluated on the SUN397[64] and Cars[31]. Figure (b) shows a contour plot of the fitted quadratic functions. Red lines represent the Pareto front in the decision variable $(c_{1},c_{2})$ space. Figure (c) shows an example of the Pareto Front. The Pareto front (Grid search) is regarded as the ground truth given the sufficient number of grid points. The Pareto front (MAP, predicted) is the amortized Pareto front. The Pareto front (MAP, real) is the Pareto front involving the same $\{(c_{1},c_{2})\}$ but re-evaluated to obtain the ground truth metrics for comparison. The yellow lines are the fine-tuned single-task models’ evaluated performance.
In this paper, we introduce a novel method that identifies the Pareto front without retraining the models to be merged. Our algorithm utilizes a quadratic approximation of the evaluation metric. Furthermore, we enhance it with a Bayesian adaptive sampling method and a nested merging scheme, which further brings down the computational cost. We validate our method across a diverse set of tasks, spanning from vision to natural language processing, and demonstrate its applicability to a variety of architectures, including ResNets [20], ViT [17], and large language models [6; 47; 55; 28]. Our results confirm that this novel approach supports the seamless integration of diverse model capabilities and aligns more closely with various real-world preference by providing a set of optimal fronts across the tasks.
Contributions
The main contributions of this paper are:
  • C1
    Design the MAP algorithm that utizes quadratic surrogate models to approximate the evaluation metric functions, which amortize the computation of Pareto fronts;
  • C2
    Propose a nested-merging MAP to decrease computational complexity from $O(N\cdot 2^{N})$ to $O(N\log_{2}^{N})$;
  • C3
    Propose Bayesian MAP, which efficiently queries the computationally expensive evaluations according to loss information.
  • C4
    To our best knowledge, this paper is the first work, that estimates the Pareto front for task-vector-based model-merging methods with low compute.
Related work
This paper is related to many existing works on multi-objective optimization, Pareto optimality, task arithmetic, federated/private learning, Bayesian methods, etc. We kindly refer the readers to Appendix A due to the page limit.
[…]
Figure 2: (a) Using the results of inference runs of merged ViT-B-32 models [45] obtained from 250 combinations of scaling coefficients, our method outperforms the direct search method in finding a set of diverse solutions for eight tasks based on the same computation expense. For both methods, we maximize a specific task while requiring all tasks to have an evaluation metric above a certain threshold ($40\%$). The bar plot for each task shows the corresponding maximized accuracy. (b) We increase the requiring bounds to ($65\%$) of the single-task model. With the same computational budget, the brute-force direct search method cannot find a feasible solution.
In Figure 2, we compare our method to direct search. We want to emphasize that, in our assumed setting, evaluation models are computationally expensive. Thus, we can only query the evaluation pipeline a limited number of times. We choose to compare to direct brute force search rather than Evolutionary algorithms like NSGA-III, etc. Evolutionary algorithms either need to query the evaluation metrics more than our computational budget or select the Pareto solutions from the pre-evaluated models. Suppose we already have a set of solutions and know their evaluation metrics, and we use evolutionary algorithms to pick the Pareto solutions among them. In that case, the set we find with the algorithms might not include all possible Pareto solution. Instead, it will likely be just a part of what we would find if we compared every possible pair of solutions directly against each other (known as brute force search). Since brute force search compares all possible solutions and finds all the Pareto optimal solutions, it’s considered a complete method. Therefore, when evaluating the effectiveness of our algorithm, it is better to compare its results to those obtained from brute force search in the case of a limited set of already-known solutions. Our code is available at https://github.com/luli-git/MAP.
2 Background
2.1 Model merging
Model merging aims to find a way to combine multiple models with the same architecture $\{\bm{\theta}_{ft}^{n}\}_{n\in[N]}$ with their initialization $\bm{\theta}_{\textup{pre}}$ in a way such that the merged model parameterized by $\bm{\theta}_{m}=\bm{\theta}_{\textup{pre}}+\sum_{n=1}^{N}c_{n}\bm{\theta}_{ft}^{n}$ perform reasonably well on all $N$ tasks [61].
A recent work by Ilharco et al. [24] introduced task arithmetic as a simple and effective way for performing model merging. The task vector for task $n$ is defined as $\mathbf{v}_{n}=\bm{\theta}_{ft}^{n}-\bm{\theta}_{\textup{pre}}$, which is the element-wise difference between the pre-trained weights and the fine-tuned weights for task $N$. To perform model merging with task vectors, we can compute $\bm{\theta}_{\textup{pre}}+\sum_{n=1}^{N}c_{n}\mathbf{v}_{n}$, where $c_{i}$ are scaling factors that have shown to be essential to the performance of the merged model [66; 67].
Denoting the metric of task $n$ as $M_{n}$, most of the existing approaches for model merging aim to improve an equal weight average metric $\frac{1}{N}\sum_{n=1}^{N}M_{n}$. This target implies that user of the algorithm has equal preferences between tasks. However, in real-world applications, users might have biased preferences for the importance of tasks, necessitating trade-offs. In such cases, the goal of model merging is no longer the equal weight average metric. Instead, we need to answer a broader question:
Given any preferences over the set of tasks, what is the best way to merge the models?
2.2 Pareto fronts
Pareto dominance
Let $X$ be a set representing the solution space, where each element $x\in X$ is a possible solution to the multi-objective optimization problem. Let there be $n$ objectives. Define an evaluation loss function $f_{i}:X\to\mathbb{R}$, where $i\in\{1,2,\cdots,n\}$. Given two solutions $x,y\in X$, we define that $x$ Pareto dominates $y$, denoted by $x\succ_{P}y$, if and only if:
$\forall i\in\{1,2,\dots,n\},f_{i}(x)\leq f_{i}(y)\text{ and }\exists j\in\{1,2,\dots,n\},f_{j}(x)<f_{j}(y)$
Pareto optimal solutions
The Pareto front is the set of solutions in the solution space $X$ that are not Pareto dominated by any other solutions in $X$.
$\text{PF}=\{x\in X\mid\not\exists y\in X\text{s.t.}y\succ_{P}x\}$ (1)
Pareto optimal solutions have been studied in multi-task (multi-objective) learning (MTL) [49; 35]. However, in most of the studies, approximating the Pareto front is computationally expensive and data inefficient. We introduce our method, MAP, a computationally efficient method to find the Pareto front for model merging.
3 Methods
3.1 Quadratic approximation of evaluation metric
Given the task vectors $\{\mathbf{v}_{n}\}_{n=1,...,N}$ and the initialization $\bm{\theta}_{\textup{pre}}\in\mathbb{R}^{d}$, we denote the merged model parameters as $\bm{\theta}_{m}(\mathbf{c})=\bm{\theta}_{\textup{pre}}+\mathbf{V}\mathbf{c}=\bm{\theta}_{\textup{pre}}+\sum_{n=1}^{N}c_{n}\mathbf{v}_{n}$, where $\mathbf{V}=\text{concat}(\mathbf{v}_{1},...,\mathbf{v}_{N})\in\mathbb{R}^{d\times N}$ is the task matrix and $\mathbf{c}=\text{concat}(c_{1},...,c_{N})\in\mathbb{R}^{N}$ is the scaling coefficients for the task vectors.
We aim to optimize the evaluation metric for all tasks, denoted by $M_{n}$, via the multi-objective optimization (MOOP) [^1] The evaluation metric $M$ can be differentiable (e.g. the mean square loss or the cross-entropy/perplexity) or not necessarily (e.g. the classification accuracy, F1 score, BLEU or Rouge):
$\min_{c_{1},\cdots,c_{N}}M_{1}(\mathbf{c}),\cdots,M_{N}(\mathbf{c})$ (2)
This problem has $N$ variables and $N$ objectives, and we seek the Pareto optimal solutions.
To do so effectively and efficiently, we use the second-order Taylor expansion to approximate $M_{n}$:
$M_{n}(\mathbf{c})$ $\equiv M_{n}(\bm{\theta}_{\text{merge}}(\mathbf{c}))=M_{n}(\bm{\theta}_{\textup{pre}})+\nabla M_{n}(\bm{\theta}_{\textup{pre}})^{\top}(\bm{\theta}_{m}(\mathbf{c})-\bm{\theta}_{\textup{pre}})$
$+\frac{1}{2}(\bm{\theta}_{m}(\mathbf{c})-\bm{\theta}_{\textup{pre}})^{\top}\mathbf{H}_{n}(\bm{\theta}_{\textup{pre}})(\bm{\theta}_{m}(\mathbf{c})-\bm{\theta}_{\textup{pre}})+R_{n}(\bm{\theta}_{m}(\mathbf{c})-\bm{\theta}_{\textup{pre}})$
$\approx M_{n}(\bm{\theta}_{\textup{pre}})+\nabla M_{n}(\bm{\theta}_{\textup{pre}})^{\top}\mathbf{V}\mathbf{c}+\frac{1}{2}(\mathbf{V}\mathbf{c})^{\top}\mathbf{H}_{n}(\bm{\theta}_{\textup{pre}})\mathbf{V}\mathbf{c}$
where $\mathbf{H}_{n}(\bm{\theta}_{\textup{pre}})=\nabla^{2}M_{n}(\bm{\theta}_{\textup{pre}})\in\mathbb{R}^{d\times d}$ is the Hessian matrix and $R_{n}(\bm{\theta}_{m}(\mathbf{c})-\bm{\theta}_{\textup{pre}})=R_{n}(\mathbf{V}\mathbf{c})$ is the third-order remainder, which is negligible when $||\mathbf{V}\mathbf{c}||^{3}=||\bm{\theta}_{m}(\mathbf{c})-\bm{\theta}_{\textup{pre}}||^{3}$ is small. Note that the second-order Taylor expansion is widely used and provides a close approximation (see Figure 2 (a) and (b) and more discussion in Appendix A), as neural networks tend to converge close to their initialization, especially for fine-tuned models: it directly motivates Newton’s method, and supports the analysis of convergence [37; 7] and the scaling laws for large models [30; 22].
Leveraging this quadratic approximation, we can define surrogate models for each task $N$, $\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c}$ and optimize the proxy problem of (2) via
$\min_{c_{1},\cdots,c_{N}}\tilde{M}_{1}(\mathbf{c}),\cdots,\tilde{M}_{N}(\mathbf{c})$ (3)
where
$\mathbf{A}_{n}=\mathbf{V}^{\top}\mathbf{H}_{n}(\bm{\theta}_{\textup{pre}})\mathbf{V}\in\mathbb{R}^{N\times N},\mathbf{b}_{n}=\mathbf{V}^{\top}\nabla M_{n}(\bm{\theta}_{\textup{pre}})\in\mathbb{R}^{N},e_{n}=M_{n}(\bm{\theta}_{\textup{pre}})+R_{n}$ (4)
Importantly, (3) is parameterized in contrast to (2), with $\frac{(N+1)(N+2)}{2}$ unknown coefficients in ($e_{n}$, $\mathbf{b}_{n}$, $\mathbf{A}_{n}$)222 1 coefficient for $e_{n}$, $N$ coefficients in $\mathbf{b}_{n}$ and $N(N+1)/2$ coefficients in $\mathbf{A}_{n}$ due to its symmetry. Thus, we can further leverage existing methods to learn the coefficients by minimizing the empirical risk over multiple $\mathbf{c}$. For instance,
$\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*}=\arg\min_{\mathbf{A}_{n},\mathbf{b}_{n},e_{n}}\sum_{\mathbf{c}\in\Omega}|M_{n}(\bm{\theta}_{\text{merge}}(\mathbf{c}))-\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})|^{2}$ (5)
where $\Omega=\{\mathbf{c}^{(1)},...,\mathbf{c}^{(K)}\}$ is the set of $\mathbf{c}$ and $M_{n}(\bm{\theta}_{\text{merge}}(\mathbf{c}))$ is the corresponding evaluation metric.
If the evaluation metric spans the whole real-number axis $\mathbb{R}$, we use the vanilla form of the surrogate model: $\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c}$. When the evaluation metric is restricted to a specific range $[l,u]$, e.g. accuracy is restricted to $[0,1]$, we would warp the quadratic part with a sigmoid function, i.e. $\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv(u-l)\sigma(e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c})+l$. Similarly, if the evaluation metric is restricted to $[l,+\infty)$, a softplus function as a wrapper would be beneficial: $\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv\text{softplus}(e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c})+l$. Please note that $u$ and $l$ mentioned here are not learnable parameters. They are specific to the feasible area of the metric.
3.2 Model merging with amortized Pareto fronts
In this section, we introduce our generalized algorithm for estimating the amortized Pareto fronts. As mentioned in Section 3.1, we approximate the evaluation metric $M_{n}(\cdot)$ by a surrogate quadratic model $\tilde{M}_{n}(\cdot)$. We then utilize $\tilde{M}_{n}(\cdot)$ to compute the amortized Pareto fronts. Please see the detailed experiments in Section 4 and the algorithm details in Algorithm 1. Different from other Pareto Multi-task learning algorithms, our algorithm
Algorithm 1 MAP
1:  Prepare models $\{\bm{\theta}_{ft}^{n}\}$ and compute task vectors $\{\mathbf{v}_{n}=\bm{\theta}_{ft}^{n}-\bm{\theta}_{\textup{pre}}\}$.
2:  for $n\in[N]$ do
3:     Sample $K$ vectors of $\mathbf{c}\in\mathbb{R}^{N}$. Denote the set as $\Omega$.
4:     for $\mathbf{c}=[c_{1},...,c_{N}]\in\Omega$ do
5:        Compute $\bm{\theta}(\mathbf{c})=\bm{\theta}_{\textup{pre}}+c_{1}\mathbf{v}_{1}+\ldots c_{N}\mathbf{v}_{N}$.
6:        Derive the evaluation metric $M_{n}(\bm{\theta}(\mathbf{c}))$.
7:     Fit the quadratic approximation surrogate model $\tilde{M}_{n}$ by learning $\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*}$ in (5).
8:  Apply MOOP algorithm (e.g. NSGA-III) to $\{\tilde{M}_{n}\}$ and get the Pareto front
3.3 Nested merging scheme
Empirically, we only need 30 pairs to get a good quality of Pareto front to merge 2 tasks, but due to the curse of dimensionality, we need exponentially more number of $(\mathbf{c},\{\tilde{M}_{n}(\theta_{m}(\mathbf{c}))\}_{n=1}^{N})$ pairs to get a good quality of Pareto front in the case of 8 tasks. Theoretically, in the best case, the points to get a good quality of Pareto front for $N$ is $O(N^{3})$. In the worst case, when the curse of dimensionality dominates, it could be $O(N\cdot 2^{N})$. Using the nested merging scheme, we reduce the computation complexity from $O(N\cdot 2^{N})$ to $O(N\log_{2}^{N})$. Please refer to Table 5 for the detailed computational complexity comparison. Algorithm details are presented in Algorithm 2.
Algorithm 2 Nested-merging MAP
0:  A predetermined preference $\textit{pref}\in\mathbb{R}^{N}$ over the $N$ tasks, the tuple of task, loss, task vector: $G_{n}=(\text{task}_{n},l_{n},\bm{\theta}_{ft}^{n})$
1:  Normalize pref to make sure the sum is 1
2:  Initialize the set $\tau=\{G_{1},\cdots,G_{N}\}$
3:  while $|\tau|>1$ do
4:     Find the pair of $(G_{i},G_{j})\in\tau$ that are closest to each other in terms of $(l_{i},l_{j})$
5:     Implement Algorithm 1 to find the Pareto front $\text{PF}_{i,j}$ between $(\tilde{M}_{i},\tilde{M}_{j})$
6:     Select $\mathbf{c}^{*}=(c_{i}^{*},c_{j}^{*})\in\mathbb{R}^{2}$ based on the Pareto front $\text{PF}_{i,j}$
7:     Merge the models by $\bm{\theta}_{\text{merge}}^{i,j}=\bm{\theta}_{\textup{pre}}+c_{i}(\bm{\theta}_{\textup{pre}}-\bm{\theta}_{ft}^{i})+c_{j}(\bm{\theta}_{\textup{pre}}-\bm{\theta}_{ft}^{j})$
8:     Calculate the weighted average loss on the two tasks $l_{ij}=\textit{pref}_{i}l_{i}+\textit{pref}_{j}l_{j}$
9:     Update $\tau$ by replacing $\{G_{i},G_{j}\}$ with $\{G_{ij}\}$, where $G_{ij}\equiv(\text{task}_{i,j},l_{ij},\bm{\theta}_{\text{merge}}^{i,j})$
10:  return  $\bm{\theta}_{\text{merge}}^{1,2,\cdots,N}$
[…]
Figure 3: (a) Discretizing of two task scaling coefficients along the angular dimension in 2D polar coordinate system (please refer to Figure 11 for an example of 3D spherical discretization.); (b) An example of nested model merging for $N=6$ models (see more details in Appendix G);
3.4 Bayesian adaptive sampling
We have discussed the strategy of bringing down the computation in the sense of big $O$ notation. It is effective when the number of tasks involved in the Pareto fronts is high.
In the case of a relatively low number of tasks ($N\leq 4$), we consider using the Bayesian adaptive sampling method inspired by Bayesian optimization.
In Algorithm 1, we only sample a single round of scaling weights $\mathbf{c}$ and query the ground truth metric by evaluation through the $M_{n}(\bm{\theta}(\mathbf{c}))$. In contrast, as for the Bayesian adaptive sampling, we sample the scaling weights $\mathbf{c}$ in multiple rounds. Each round, the sampling strategy depends on the previous sampling $\mathbf{c}$ and its evaluation metrics. The process begins with an initial uniform sampling of scaling coefficients, $\{(c_{1},\cdots,c_{N})_{i}\}_{i=0}^{n_{0}}$, from $[0,1]^{N}$. For simplicity, we define $\mathbf{c}_{i}$ as $(c_{1},\cdots,c_{N})_{i}$. We evaluate the merged model $\theta_{m}(\mathbf{c}_{i})$ across tasks $1$ to $N$ and compute the corresponding $L2$ loss. We then discretize the regions of $\mathbf{c}$ into joint bins within a hyper-spherical coordinate system and discretize along the angular dimensions. Please refer to figure 11 as examples in 2-dimensional and 3-dimensional discretization. The posterior distribution is computed proportional to an acquisition function (e.g., upper confidence bound) of the approximation loss within specific bins.
We iteratively update each surrogate model for task $n$. See the algorithm 3 for a simplified version of the process. For more details, please refer to the LABEL:{alg:BASM} in the appendix. After meeting the stopping criterion, we use the surrogate models for tasks $\{t\}_{n=1}^{N}$ to generate a lot of $(\mathbf{c},\{\tilde{M}_{n}(\theta_{m}(\mathbf{c}))\}_{n=1}^{N})$ and apply MOOP algorithm (e.g. NSGA-III) to $\{\tilde{M}_{n=1}^{N}\}$ to calculate the Pareto front.
Algorithm 3 Bayesian MAP
0:  Number of iterations $J$, Buffer $\mathcal{B}$, Pretrained model $\bm{\theta}_{\textup{pre}}$, Task vectors $\mathbf{V}=(\mathbf{v}_{1},\cdots,\mathbf{v}_{n})$, Evaluation function $M_{n}(\cdot)$ for task $n$.
1:  Initialize $\mathcal{B}$ to an empty set.
2:  for each iteration $j=0$ to $J$ do
3:     Sample scaling coefficients $\{\mathbf{c}_{i}\}_{i=1}^{n_{j}}$ based on updated distribution from previous results (uniformly sampled if $j=0$).
4:     for each scaling coefficient $i=1$ to $n_{j}$ do
5:        Adjust the model with $\theta_{m}(\mathbf{c}_{i})=\bm{\theta}_{\textup{pre}}+\mathbf{c}_{i}\cdot\mathbf{V}$.
6:        Evaluate this model configuration $m_{n,i}=M_{n}(\theta_{m}(\mathbf{c}_{i}))$ and store $(\mathbf{c}_{i},m_{n,i})$ in $\mathcal{B}$.
7:     Update the surrogate model based on the $(\mathbf{c},\{\tilde{M}_{n}(\theta_{m}(\mathbf{c}))\}_{n=1}^{N})$ pairs in $\mathcal{B}$.
8:     Estimate acquisition function (UCB) for different model configurations.
9:     Adjust the sampling strategy for the next iteration based on these evaluation metrics.
10:  return
4 Experiments
In this section, we present our empirical findings across a wide range of tasks including zero-shot classification (CLIP-based) and LLMs (Llama3-based).
4.1 Experiment setup
Table 1: Experiment setup in terms of tasks and models
Task type Metric # of total tasks Model type
Zero-shot Classification (normal) Accuracy 8 ViT-B/32 (CLIP)
Zero-shot Classification (medical) Accuracy 2 ViT-B/32 (CLIP)
Language Generation Loss/Perplexity 4 Llama3-8B
Image Classification Accuracy 3 ResNet-18
Datasets and models
We study multi-task model merging on eight normal zero-shot image classification datasets following [24]: SUN397 [64], Cars [31], GTSRB [51], MNIST [33], EuroSAT [21], SHVN [40], DTD [11], RESISC45 [9]. We use the ViT-B/32 architecture in CLIP [45] as the pre-trained model for the experiments on vision tasks in the main text. We show the result of them in the main pages. For the rest of the experiments, due to the page limits, we show them in the appendix. The results of the datasets and tasks we experimented on are as follows: a zero-shot medical chest X-ray image classification task to show our model merging scheme also works in a real-world application domain [58]; 4 fine-tuned Llama3 models in different languages: French, Arabic, Chinese, and Japanese  [^3] All the fine-tuned language Llama3 models can be found on Huggingface. The IDs of the models are: French: jpacifico/French-Alpaca-Llama3-8B-Instruct-v1.0; Arabic: MohamedRashad/Arabic-Orpo-Llama-3-8B-Instruct; Chinese: shenzhi-wang/Llama3-8B-Chinese-Chat; Japanese: haqishen/Llama-3-8B-Japanese-Instruct.; 3 vision tasks on the ResNet18 architecture: CIFAR-10 [32], Flowers-102 [41], and GTSRB [51].
4.2 Baseline and metrics
Baselines
Our baseline method for obtaining Pareto fronts is the brute-force direct search. When the number of tasks is low ($N<4$), we can sample enough grid points of $\mathbf{c}$ and query the corresponding evaluation metrics $\{M_{n}(\theta(\mathbf{c}))\}_{n=1}^{N}$. We can then directly use the resulting evaluation metrics to find the Pareto front by direct comparisons of each task performance of the merged model by $c_{i}\in\mathbf{c}$. When the number of tasks is low, we can regard the resulting Pareto front as the ground truth Pareto front. However, when the number of tasks grows, the required number of $(\mathbf{c},\{M_{n}(\theta(\mathbf{c}))\}_{n=1}^{N})$ pairs grows exponentially, which is way larger than the points we evaluated. Thus, when the number of tasks is high ($N\geq 4$), the results from the brute-force direct search can no longer be considered as ground truth.
In addition to the brute-force method, we compare with other model merging methods including SLERP [50], TIES-merging [66], Task Arithmetic [24] with a single scalar, and Task Arithmetic with preferences as scalars, DARE combined with Task Arithmetic [68], and DARE combined with TIES-merging.
Win rate
We used the win rate to measure how often the Pareto front found by MAP outperforms the Pareto front found by the baseline in terms of multiple objectives. Let $PF_{MAP}$ and $PF_{baseline}$ represent the set of solutions in the Pareto fronts obtained from the MAP and the baseline methods, respectively. Each solution in these sets is a vector in $\mathbb{R}^{N}$, where $N$ is the number of objectives or tasks. We sampled $K=\min(100,|PF_{baseline}|)$ points from each of the two Pareto fronts, denoted as $\mathbf{c}_{k}^{MAP}$ and $\mathbf{c}_{k}^{baseline},k=1,\ldots,K$. Then, we compared $M_{n}(\theta(\mathbf{c}_{k}^{MAP}))$ and $M_{n}(\theta(\mathbf{c}_{k}^{baseline}))$ pairwise for $k=1,\ldots,K$ and $n=1,\ldots,N,$ resulting in $K\times K\times N$ comparisons. The ratio of instances where $M_{n}(\theta(\mathbf{c}_{k}^{MAP}))>M_{n}(\theta(\mathbf{c}_{k}^{baseline}))$ is computed as the win rate of our amortized Pareto front $PF_{MAP}$.
4.3 Win rate of MAP over the direct search method
Table 2 shows the results for the win rate. In conclusion, when the number of tasks ($N<4$), we can use only query 30 and 50 scaling coefficients to get similar performance with the ground truth Pareto front. For the high number of task scenarios ($N\geq 4$), MAP performs much better than the brute-force (not ground truth) Pareto solutions.
Table 2: Win rate of the amortized Pareto front and the Pareto front by brute-force direct search. The number of points is the number of points each algorithm used to find the Pareto front. Note that the number of evaluations by each method is the number of points multiplied by the number of tasks. The number of points per dimension is the number of points raised to the power of $\frac{1}{N}$, which measures the sparsity of points the direct search method uses per dimension to approximate the Pareto front.
# of tasks # of pts direct search # of pts per dim # of pts (MAP) Win rate (MAP) $R^{2}$ (MAP)
2 200 14.14 30 49.81% (±$(\pm 0.30)$ 0.953 (±$(\pm 0.018)$
3 300 6.69 50 46.90% (±$(\pm 0.71)$ 0.980 (±$(\pm 0.003)$
4 300 4.16 60 50.67% (±$(\pm 2.44)$ 0.984 (±$(\pm 0.004)$
5 500 3.47 85 53.00% (±$(\pm 1.88)$ 0.941 (±$(\pm 0.019)$
6 500 2.82 100 60.71% (±$(\pm 1.34)$ 0.941 (±$(\pm 0.030)$
7 1000 2.68 140 63.42% (±$(\pm 1.91)$ 0.891 (±$(\pm 0.024)$
8 1000 2.37 250 65.58% (±$(\pm 0.94)$ 0.868 (±$(\pm 0.028)$
4.4 MAP offers a well-distributed set of Pareto optimal solutions compared with other baseline methods
We compared the performance of MAP with TIES-merging [65], TIES-merging with DARE [68], Task Arithmetic with DARE, Task Arithmetic with normalized preference as scaling coefficients [25], SLERP [50]. For dimension 2, we can directly visualize the results, as shown in Figure 4. In addition, we show that MAP is a plug-in that can be directly combined with other task vector-based model merging methods where the users can control the preferences using some scalars. In Figure 4, we show the Pareto front found by MAP-DARE.
Refer to captionRefer to captionRefer to caption
Figure 4: The Pareto fronts obtained using MAP with Task Arithmetic, MAP with Task Arithmetic and DARE. We sampled 10 Pareto solutions from the predicted front by MAP and evaluated them to obtain the real values. We plotted the results obtained using TIES-merging, Task Arithmetic (TA) with a single scalar for all tasks, Task Arithmetic with preferences as scalars, TA combined with DARE (DARE-TA), TIES-merging combined with DARE (DARE-TIES), and SLERP.
We can see that none of the baseline methods can directly dominate all Pareto solutions found by MAP, and most of them lie either within the Pareto front found by MAP. In addition, we sampled 10 points from the predicted Pareto front by MAP and evaluated them (MAP Pareto solutions, real) to confirm that they indeed lie close to the predicted Pareto front. This evidence further confirms the usefulness of MAP as a general strategy for finding a set of diverse and well-distributed Pareto solutions that none of the existing model merging methods can substitute.
For dimensions higher than 2, we use two metrics to compare the Pareto front found by MAP with other baseline methods. We sample 20 preference vectors and normalize them. Then, we pick the solution by MAP that maximizes the preference-weighted sum of accuracies and compare it with those by baseline methods. The results are shown in Table 3. We also compute the win rate, which is the proportion of preferences where the preference-weighted sum of accuracies of MAP is higher than that of all the baseline methods. We can observe that even in higher dimensions, when the user has certain preferences, MAP is still useful and can better accommodate user preferences. Please note that when using MAP, we do not require the user to pre-specify their preference vector. This preference vector based evaluation is only for evaluation purposes.
Table 3: We compared MAP with a set of baseline methods by sampling a set of 20 normalized preference vectors and computing the preference-weighted sum of accuracies and win rate. The win rate is defined as the proportion of the 20 preference vectors where the preference-weighted sum of accuracies of MAP is higher than those of the baseline methods. $\uparrow$ indicates higher is better. The number after $\pm$ is the standard deviation. Please refer to Table 2 for the number of scaling coefficient vectors used for different number of tasks.
Preference weighted sum of accuracies ($\uparrow$)
# tasks 2 3 4 5 6 7 8
Single task models 75.84$\pm$1.76 77.03$\pm$1.84 82.43$\pm$4.40 87.69$\pm$4.50 88.52$\pm$4.02 89.26$\pm$3.58 90.62$\pm$2.52
MTL 73.63$\pm$0.30 75.13$\pm$1.00 80.10$\pm$2.79 84.93$\pm$3.58 86.78$\pm$2.94 87.40$\pm$2.56 89.11$\pm$2.36
MAP 70.70$\pm$0.21 69.05$\pm$1.41 72.84$\pm$1.05 77.31$\pm$0.83 74.26$\pm$0.52 73.40$\pm$-0.14 72.96$\pm$0.73
Model soups 67.79 $\pm$ 1.46 64.25 $\pm$ 2.15 66.04 $\pm$ 3.22 67.01 $\pm$ 3.42 63.11 $\pm$ 1.99 63.35 $\pm$ 2.17 64.36 $\pm$ 2.77
TIES-merging 69.30$\pm$0.33 67.60$\pm$0.58 71.79$\pm$2.93 76.49$\pm$3.10 73.74$\pm$2.96 72.54$\pm$2.87 72.24$\pm$1.91
DARE-TIES 67.62$\pm$1.65 66.49$\pm$2.34 71.39$\pm$4.45 74.55$\pm$4.55 73.34$\pm$4.10 71.43$\pm$3.84 71.89$\pm$2.86
Task Arithmetic 70.73$\pm$1.84 61.15$\pm$2.33 52.69$\pm$4.23 61.58$\pm$4.62 51.37$\pm$3.84 39.79$\pm$3.97 60.77$\pm$2.84
TA with preference as weights 69.22 $\pm$ 1.4 66.88 $\pm$ 2.37 68.73 $\pm$ 5.48 71.92 $\pm$ 5.5 68.13 $\pm$ 4.69 68.14 $\pm$ 4.2 68.17 $\pm$ 2.89
DARE-TA 70.61$\pm$0.22 64.18$\pm$1.24 58.04$\pm$8.19 65.39$\pm$7.03 56.76$\pm$7.01 46.75$\pm$5.73 64.51$\pm$3.81
Win rate of MAP over baseline methods ($\uparrow$)
# tasks 2 3 4 5 6 7 8
MAP vs TIES-merging 75$\pm$9.80 60$\pm$10.93 70$\pm$10.49 75$\pm$9.39 65$\pm$10.85 65$\pm$11.66 70$\pm$10.25
MAP vs DARE-TIES 90$\pm$6.81 80$\pm$9.11 75$\pm$9.87 90$\pm$6.70 65$\pm$10.59 60$\pm$11.38 75$\pm$9.70
MAP vs TA 65$\pm$10.91 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0
MAP vs DARE-TA 65$\pm$10.64 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0
MAP vs TA with preference 85.0 $\pm$ 0.08 85.0 $\pm$ 0.08 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0
MAP vs Model soup 90.0 $\pm$ 0.07 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0 100.0 $\pm$ 0.0
4.5 Results on using nested merging to merge 8 models
Table 4: We compared nested-merging MAP with MAP and baseline methods in merging 8 models using a set of 10 preference vectors in terms of preference-weighted sum of accuracies.
Metric Single-task models MAP-nested MAP-plain TIES-merging DARE-TIES Task Arithmetic (TA) DARE-TA
Preference weighted sum ($\uparrow$) 90.05 $\pm$ 3.02 67.05 $\pm$ 3.89 72.12 $\pm$ 3.78 70.79 $\pm$ 3.81 70.51 $\pm$ 3.58 59.44 $\pm$ 5.68 63.14 $\pm$ 4.91
We further evaluate the effectiveness of nested merging MAP. Supplemental Figure 6 shows the intermediate Pareto fronts obtained using nested merging MAP sequentially. In addition, we compared the performance of nested merging MAP (Algorithm 2) with MAP (Algorithm 1). The results are shown in Table 4. The results confirm that nested-merging MAP can obtain comparable results to those obtained by MAP while using much fewer evaluation runs. Specifically, for nested merging MAP (check  Figure 6 for the process), we use 20 scaling coefficients for each pairwise merge. Thus, in the first round of merging, the number of evaluations is $20\times 2~{}\text{tasks}=40~{}\text{times}$. In the second round of merging, the number of evaluations is $20\times 4~{}\text{tasks}=80~{}\text{times}$. In the last round of merging, the number of evaluations is $20\times 8~{}\text{tasks}=160~{}\text{times}$. According to  Figure 6, there are 4 first rounds in the first merging, 2 second rounds in the second merging and 1 last round. Thus, in total, we run $40\times 4+80\times 2+160\times 1=480$ evaluations. In contrast, when running the vanilla MAP (Algorithm 1), we used 280 coefficients and evaluate $280\times 8=2240~{}\text{times}$. In addition, while being suboptimal to MAP (Algorithm 1), nested-merging MAP can still outperform other baseline methods such as TA and DARE-TA in accommodating various user preferences with relatively low computational cost, especially when the number of tasks is high.
4.6 Bayesian MAP experiments
Please refer to Appendix D for Bayesian MAP experiments.
5 Conclusion and Limitations
Our method (MAP) can serve as an out-of-the-box plug-in that is very easy to combine directly with other task-vector based model-merging methods, such as tasks arithmetic [24] and DARE [68], etc. From Section 3 and Section 4, we have shown that our methods Algorithm 1, algorithm 2, and Algorithm 3 can efficiently reduce the computational cost from $O(N2^{N})$ to $O(N\log_{2}{N})$. However, we would like to point out some limitations of our method. First, during the design of Algorithm 1, we do not handle the situation of a non-convex Pareto front. In the case of two tasks, it is likely that the Pareto fronts with proper trade-offs would have a convex Pareto front. However, if one deals with more tasks, the ground truth Pareto fronts are likely non-convex, and Algorithm 1 ends up approximating a convex hull. A quick but naive solution would be to implement Algorithm 2 in this case, so that the Pareto fronts between two tasks would be more likely to be convex. Second, we do not have a detector algorithm to detect if two tasks have proper trade-offs. Approximating a Pareto "front" with only a single Pareto solution likely results in very unstable and poorly-performed results. In this case, practitioners should be informed.
Appendix A Related Work
Multi-objective optimization
Multi-objective optimization (MOOP) aims at identifying a variety of Pareto solutions, each offering different trade-offs, instead of just a single solution. In fact, the multi-task problem is approached from a MOOP view point [49; 35], which leverages a rich literature of algorithms including MGDA, IMTL, GradNorm, RLW, PCGrad, scalarization, and so on. We note that these training algorithms iteratively optimize over the $\mathbb{R}^{d}$ model parameters and can be computationally costly (in order to instantiate and modify the per-task gradient). In contrast, our algorithm is a post-training MOOP over the $\mathbb{R}^{N}$ scaling coefficients, which is significantly low-compute.
Many methods have been proposed to solve MOOP and derive the Pareto optimal solutions. One may leverage the scalarization technique to transform a multi-objective problem into many single-objective ones, aiming to find one solution for each problem and approximate the Pareto front. Additionally, one can solve the Karush-Kuhn-Tucker conditions, which are easy to work with, given that the parameters have a closed-form solution under our quadratic approximation.
To select an appropriate MOOP algorithm, we need to take into account the computational costs and the quality of the Pareto front. For instance, many indicator-based and selection-based methods (e.g. the hypervolume indicator [3]) have the time complexity that grows super-polynomially with the number of objectives. Indeed, for $N\geq 3$, MOOP is termed as many-objective optimization, which is significantly challenging for well-established algorithms (e.g. NSGA-II and SPEA2) and requires more advanced algorithms such as $\epsilon$-MOEA [16], MSOPS [23], SMS-EMOA [5], and NSGA-III [15], or the KKT approach [4].
Task arithmetic
Task arithmetic, a method of model merging that has drawn increasing attention, applies a weighted average of models (controlled by the scaling coefficients) to obtain high performance on multiple tasks simultaneously. Several existing works have studied ways to select the scaling coefficients [24; 66; 67]: [24] simply uses equal scaling, similar to the Model Soup approach, which is heuristic and suboptimal, plus the applicability is limited to certain tasks and small scales; [67] aims to learn the optimal scaling weights by using the Shannon entropy as a surrogate objective function to the cross entropy loss. However, it requires the use of unlabeled test data from all tasks. In most real-world applications, data cannot be directly shared, which is what makes model merging appealing initially [36]. In addition, such an approach only works for classification tasks due to the use of cross-entropy loss.
[62] proposed ‘WiSE-FIT’ to perform robust fine-tuning by computing the weighted average of the pre-trained model parameters with the parameters of the models fine-tuned on different tasks, with different weights determining different trade-offs between pre-training and fine-tuning task performance. Task Arithmetic [24] adds the weighted sum of the differences between the fine-tuned model and the pre-trained model to the weights of the pre-trained model. Again, the weights are used to distinguish the importance of various models. In addition, [43] explores how fine-tuning models in their tangent space enhances weight disentanglement, leading to improved performance in task arithmetic for editing pre-trained vision-language models.
Different from task arithmetic based model merging approaches, Ainsworth et al. [1] show that even if the models are not finetuned from the same pretrained model, practitioners could still merge the models by permutations of the rows and columns of the weight matrices. Similarly, Jin et al. [29] also introduce a dataless knowledge fusion method which is not based on task vectors that merges fine-tuned language models in parameter space without requiring access to their original training data, outperforming existing baselines and offering an efficient alternative to multi-task learning. Daheim et al. [14] proposes an uncertainty-based gradient matching method for model merging that improves performance and robustness by reducing gradient mismatches between models trained on different datasets.
Second-order Taylor expansion
In deep learning, the second-order Taylor expansion and thus the quadratic approximation on the evaluation metric is an important technique, that characterizes the metric landscape. For example, the Newton-Ralphson method is derived from it: $\mathbf{w}_{t}-\mathbf{w}_{t+1}=\mathbf{H}^{-1}\mathbf{G}=\text{argmin}_{\mathbf{v}}M(\mathbf{w}_{t}-\mathbf{v})$ given that $M(\mathbf{w}_{t}-\mathbf{v})=M(\mathbf{w}_{t})-\mathbf{v}\mathbf{G}+\frac{1}{2}\mathbf{v}^{\top}\mathbf{H}\mathbf{v}$. The quadratic approximation is also combined with Lipschitz smoothness ($L$) in the classic convergence analysis of SGD [8; 19], through $M(\mathbf{w}_{t}-\eta\mathbf{G})\leq M(\mathbf{w}_{t})-\eta||\mathbf{G}||^{2}+\frac{L\eta^{2}}{2}||\mathbf{G}||^{2}$. Interestingly, although the approximation is only accurate in the local neighborhood, it can be used to indicate the global convergence over the iterations. One reason is that state-of-the-art neural networks are usually over-parameterized and undergo lazy training, meaning that the converging parameters are close to the initialization [10; 18; 2]. Thus, the local approximation informs the global convergence behavior. In particular, this approximation has played important roles in the scaling laws of neural networks (e.g. Equation 3.3 in [52]) that predict the performance and help select hyperparameters before training actually takes place.
Federated/Private Learning
Because the model merging is a post-training approach, it can naturally be used in federated learning, which protects both the model privacy and data privacy. To be specific, for data privacy, the only operation that requires the data is line 6 in Algorithm 1, where each dataset can be kept at one site without sharing with other sites: the $i$-th site holding its own data will take in the model $\bm{\theta}(\mathbf{c})$ and only share $M_{n}(\bm{\theta}(\mathbf{c}))\in\mathbb{R}$; for model privacy, the only operator that requires an aggregation is line 5 in Algorithm 1, where each model can be kept at one site and the aggregation is privacy-preserving using a safe center site, or the secure multi-party computation without a center site.
Pareto fronts for Multi-task Learning
Recent advancements in multi-task learning (MTL) have seen diverse approaches to Pareto Front learning in machine learning.
Lin et al. [35] developed a constrained optimization approach to find multiple Pareto optimal solutions representing different trade-offs among tasks in multi-task learning and solving decomposed subproblems in parallel using gradient-based optimization. It requires separate training runs for each solution. Navon et al. [39] and Lin et al. [34] employed hypernetworks for continuous Pareto Front approximation, generating target network weights based on preferred trade-offs. However, the size of these hypernetworks can become prohibitively large and need to be properly trained as well. Ruchte and Grabocka [48] introduced a memory-efficient method by augmenting network inputs with desired trade-offs, although this may obscure the network’s conditioning due to nonlinear dynamics, and it is also based on the gradient updating method.
Model-merging Applications in LLM
In the realm of model merging applications for language model preferences, recent research has made significant progress.Rame et al. [46] and Jang et al. [27] introduced "rewarded soups" and "personalized soups", which utilize model soup to interpolate weights of networks fine-tuned on diverse rewards to achieve Pareto-optimal generalization across preference spaces and address the challenge of aligning large language models with individual perspectives. Zhong et al. [70] developed "Panacea", which reframes alignment as a multi-dimensional preference optimization problem, using singular value decomposition-based low-rank adaptation to guide model behavior with a low-dimensional preference vector. To address the limitations of scalar rewards in RLHF, Zhou et al. [71] introduced Multi-Objective Direct Preference Optimization, an RL-free algorithm that extends Direct Preference Optimization to handle multiple alignment objectives efficiently. Finally, Wang et al. [57] proposed the Directional Preference Alignment framework, which incorporates multi-objective reward modeling to represent user preferences, offering intuitive arithmetic control over LLM generation for diverse user preferences.
Bayesian Optimization
Bayesian optimization has been widely used in scenarios that require efficient exploration of parameter spaces, particularly when evaluating the performance of each configuration is costly or time-consuming. This method is especially advantageous in machine learning and hyperparameter tuning, where traditional optimization techniques may be computationally prohibitive. Popular Bayesian optimization methods and applications are [59; 26; 38; 44; 53].
Appendix B Performance of the quadratic approximation
We further validated the quadratic approximation of the surrogate models and the true performance by measuring the out-of-sample $R^{2}$. Figure 5 shows the relationship between the average $R^{2}$ across all tasks and the number of total points used to fit the quadratic function. As we can see from Figure 5 (a), the average $R^{2}$ values are close to $1$ when the dimension of the decision space is low ($N<4$) when using around $30$ points. However, as the dimension of the decision space increases ($N\geq 4$), the average $R^{2}$ depends more heavily on the number of points (Figure 5 (b)).
[…]
Figure 5: Average $R^{2}$ value as a measure of the quality of fitting the surrogate model for merging different numbers of tasks. (a) Number of tasks is 2 and 3; (b) Number of tasks is 4, 5, 6, and 7.
Appendix C Nested-Merging MAP
C.1 Time Complexity of Nested-Merging MAP
Table 5: Computational cost of model merging for $N$ models.
# evals per task minimum # evals (total) # evals (total)
Naive model merging $O(N^{2})$ $O(N^{3})$ $O(N\cdot 2^{N})$
Nested model merging $O(1)$ $O(N\log_{2}^{N})$ $O(N\log_{2}^{N})$
To estimate the computational complexity of nested-merging MAP, we denote N as the number of tasks. The number of total evaluations needed for nested-merging MAP is: $N/2\times 2+\cdots+N/{2^{m}}\times 2^{m}=O(N\log_{2}^{N})$ where $2^{m-1}<N\leq 2^{m}$.
Detailed calculations are as follows. When the number of tasks is 8, estimating the surrogate model in MAP (Algorithm 1) with 8 degrees of freedom in scaling coefficient vectors could still be costly. In practice, we need to sample about 300 sets of scaling coefficient vectors, get their corresponding merged models and thus take $300\times 8=2400$ times of evaluation to implement MAP (Algorithm 1). However, if we implement Nested-Merging (NMMAP) (Algorithm 2), we only need to evaluate 20 merged models for each round of 2-task merging. In the first rounds, there are 4 2-task mergings running in parallel, each of them evaluates 2 tasks. Thus, it takes $4\times 20\times 2=160$ times of evaluation. In the second round, there are 2 2-task mergings running in parallel, each of them evaluates 4 tasks. It is 4 tasks rather than 2 tasks because when we merge the models: model for task a, b ($model_{a,b}$) and model for task c, d ($model_{c,d}$), we need to evaluate the merged model: $model_{a,b,c,d}$ on Task a, b, c, d. Thus, it takes $2\times 20\times 4=160$ times of evaluation. In the last round, there is only one 2-task merging and evaluation on 8 tasks. It takes $1\times 20\times 8=160$ times of evaluation. In total, NMMAP takes 480 times of evaluations, which is far less than 2400 evaluations.
In conclusion, we can see the number of rounds can be calculated by $\log_{2}(N)$, in each round, we need to evaluate $T\cdot N/2^{i}\cdot 2^{i}=TN$ where $T$ is the number of scaling coefficient vectors needed to be evaluated when the number of tasks is 2. In the above example, $T=20$. Thus, the time complexity is $O(TN\log_{2}(N))$. We can rewrite it to $O(N\log_{2}(N))$ if ignoring $T$.
C.2 Pareto Fronts Obtained Using Nested Merging
In our experiments, we explored the strategy of decomposing tasks into pairs, identifying the Pareto front for each pair, and then sequentially merging these binary-combined models two at a time. Our goal was to determine if this incremental merging approach affects performance negatively when compared to merging multiple models in a single operation.In Figure 6, we show the intermediate Pareto fronts obtained when merging 8 models using nested-merging MAP, where we merge two models at a time. The figures show the intermediate Pareto fronts obtained, where $A\_B$ means the model obtained by merging model A and model B.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Illustration of the sequential steps involved in merging 8 models using nested-merging MAP (left to right, top to bottom). The figures show the intermediate Pareto fronts obtained, where $A\_B$ means the model obtained by merging model A and model B.
C.3 Nested merging with rule-based preferences
One of the main advantages of MAP is that a numeric preference vector is not required. In addition to the preference-based evaluation we performed in Table 4, we performed another set of experiments merging four models using rule-based performance. Rule-based preference means preferences such as "I want the merged model to have performance above 70% for tasks A, B, and C, while maximizing the performance on task D". Such preferences can be easily accommodated using MAP or Bayesian MAP, where the complete Pareto front can be obtained. However, for nested-merging MAP, since we merge two models at a time, obtaining the full Pareto front at one time is not possible. Thus, we conducted experiments on merging four models, where the merged model for two models is not picked by numeric preference vectors, but some rule-based preference.
We performed nested model merging on four tasks and compared its performance with merging four tasks directly. When merging two models (SUN397 and DTD, GTSRB and Cars), we used rule-based preferences, which requires all models to achieve 70% of the single-task fine-tuned model, while maximizing the performance on the first task (SUN397 (1) and GTSRB (3)).
Table 6: Accuracy on four tasks when using nested merging vs direct merging. We first merge models 1 and 2, and models 3 and 4 in parallel, using the preference that requires all tasks to have a normalized performance of at least 70% while optimizing the performance on tasks (1 and 3). We then compute the Pareto front for the two combined models. The table shows the highest accuracies on each task, while requiring all tasks to achieve a normalized performance of 65%.
SUN397 (1) DTD (2) GTSRB (3) Cars (4)
Nested merging 73.24 (±$(\pm 1.03)$ 50.69 (±$(\pm 1.67)$ 96.26 ±$\pm(0.31)$ 56.20 (±$(\pm 3.46)$
Directly merge 4 models 68.36 (±$(\pm 2.63)$ 61.70(±$(\pm 6.78)$ 98.97 (±$(\pm 0.41)$ 68.24(±$(\pm 3.38)$
Appendix D Bayesian MAP
Generational distance and inverted generational distance
We evaluated the quality of the Pareto front in capturing the shape of the ground truth Pareto front by measuring how much the predicted Pareto front converges to the ground truth Pareto front by calculating the generational distance (GD) [56] and how much the predicted Pareto front covers the ground truth Pareto front by calculating the inverted generational distance (IGD) [12]. GD and IGD are standard measures used in evolutionary multi-objective optimization to evaluate the solutions found by the evolutionary algorithms.Given two solution sets $PF_{i}=\{\tilde{M}_{1}^{i}(\theta_{m}(\mathbf{c})),\ldots,\tilde{M}_{N}^{i}(\theta_{m}(\mathbf{c}))\},i=1,2$, the GD and IGD metrics are defined as
$GD(PF_{1})\equiv\frac{1}{K}\left(\sum_{i=1}^{K}d_{i}^{p}\right)^{1/p}\text{ and }IGD(PF_{1})\equiv\frac{1}{M}\left(\sum_{i=1}^{M}\tilde{d}_{i}^{p}\right)^{1/p}$ where $d_{i}$ is the minimal Euclidean distance from $\tilde{M}_{1}^{1}(\theta_{m}(\mathbf{c}))$ to $PF_{2}$ and $\tilde{d}_{i}$ is the minimal Euclidean distance from $\tilde{M}_{1}^{2}(\theta_{m}(\mathbf{c}))$ to $PF_{1}$.
Performance of Bayesian MAP
We further improve the efficiency of MAP by proposing an adaptive Bayesian sampling algorithm. This Bayesian approach samples the points from regions with the highest level of uncertainty, where uncertainty is quantified with the Upper confidence bound (UCB). Please refer to Algorithm 4 for more details about the algorithm.
Please refer to Appendix D as the experiment result comparing Bayesian MAP Algorithm 4 with MAP Algorithm 1. The very left column shows the name of the 2 tasks being merged. Below, we define points (pts) as scaling coefficients and evaluation metrics of the corresponding merged models. Please note that the result shown in this figure is the mean of 7 merging tasks that merge 2 models: DTD+Cars, DTD+RESISC45, EuroSAT+RESISC45, GTSRB+Cars, GTSRB+RESISC45, RESISC45+SVHN, SUN397+SVHN. Please also check the Pareto front estimated by Bayesian MAP with only one iteration (6+3 pts) in Figure 7.
The experiments are initialized with 6 pairs of $\mathbf{c},\{M_{n}\}_{n=1}^{N}$ (iter 0). In every following iteration, we sample more points following the Bayesian adaptive sampling algorithm. We compare the Bayesian adaptive sampling beginning with 6 points and adding 3 additional points (6+3 pts) with running Algorithm 1 with 9 points in a row. We also compare the Bayesian adaptive sampling beginning with 6 points and adding 3 additional points for 2 times (6+2 $\times$ 3 pts) with running Algorithm 1 with 12 points in a row. We show that, under most cases, utilizing the same number of points, Bayesian adaptive sampling performs better than the run-in-a-row scheme in Algorithm 1.
In conclusion, when the number of data points (scaling coefficients and evaluation metrics of the corresponding merged models) is small, the Bayesian MAP Algorithm 4 performs better than MAP Algorithm 1. As the number of data points increases, their performance becomes closer. Thus, we recommend implementing Bayesian MAP when the computational resource is very limited and the number of models (tasks) to merge is not high.
captionBayesian MAP compared to MAP. The x-axis represents the number of points used by either MAP or Bayesian MAP, while the y-axis represents the value for IGD or GD. We compared MAP with 6, 9, and 12 points, and Bayesian MAP with 6 initial points, sampling 3 more points each round for two rounds.
[…]
Figure 7: (a) This plot shows a failure case of the amortized Pareto front when we only have 6 initial randomly sampled pairs of $(\mathbf{c},\{M_{i}(\theta_{m}(\mathbf{c}))\}_{i=1}^{N})$ (b) After one iteration of Bayesian adaptive sampling for 3 more pairs (9 in total), the amortized Pareto front is much better than the initial Pareto front.
Appendix E Proof: Negligibility of the Remainder in Multivariate Taylor Series
Corollary: Accurate Local Approximation by Taylor Expansion
If $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $(k+1)$ times continuously differentiable in a neighborhood around a point $a\in\mathbb{R}^{n}$, then the Taylor polynomial $T_{k}(x)$ of order $k$ provides an accurate approximation of $f(x)$ when $x$ is sufficiently close to $a$. Furthermore, the remainder term $R_{k}(x)$ becomes negligibly small as $||x-a||$ approaches zero, assuming that the $(k+1)$th derivatives of $f$ are bounded by a constant $M$ in the neighborhood between $a$ and $x$.
Proof
Consider the Taylor series expansion of $f$ around the point $a$, truncated at order $k$:
$T_{k}(x)=\sum_{|\alpha|\leq k}\frac{D^{\alpha}f(a)}{\alpha!}(x-a)^{\alpha}$
where $\alpha$ is a multi-index of non-negative integers, $D^{\alpha}f(a)$ denotes the partial derivatives of $f$ at $a$ corresponding to $\alpha$, and $(x-a)^{\alpha}=(x_{1}-a_{1})^{\alpha_{1}}\cdots(x_{n}-a_{n})^{\alpha_{n}}$.
Assumptions
  1. 1.
[…]
  2. 2.
    Bounded Derivatives: There exists a constant $M$ such that for all multi-indices $\alpha$ with $|\alpha|\equiv k+1$, the norm of the tensor $D^{\alpha}f$ evaluated at any point $\xi$ between $a$ and $x$ is bounded by $M$.
$\|D^{\alpha}f(\xi)\|=\sup_{\|v_{1}\|=1,\ldots,\|v_{k+1}\|=1}|D^{\alpha}f(\xi)(v_{1},\ldots,v_{k+1})|\leq M$
The remainder term of the Taylor series expansion is given by:
$R_{k}(x)=\sum_{|\alpha|=k+1}\frac{D^{\alpha}f(\xi)}{\alpha!}(x-a)^{\alpha}$
Given the assumptions, we estimate:
$|R_{k}(x)|\leq\sum_{|\alpha|=k+1}\frac{\|D^{\alpha}f(\xi)\|}{\alpha!}\|x-a\|^{k+1}\leq\sum_{|\alpha|=k+1}\frac{M}{\alpha!}\|x-a\|^{k+1}$
As $\|x-a\|\rightarrow 0$, the term $\|x-a\|^{k+1}$ goes to zero. Thus, the remainder term $R_{k}(x)$ becomes arbitrarily small, making it negligible.
In conclusion, under the stated assumptions, the Taylor series truncated at order $k$, $T_{k}(x)$, provides an accurate approximation of $f(x)$ near $a$, and the remainder $R_{k}(x)$ can be ignored as $\|x-a\|\rightarrow 0$ and the higher-order derivatives remain bounded by $M$.
Appendix F Additional Experiment Results
F.1 Zero-shot Medical Image Classification
In addition to natural images, we used another dataset consisting of over 112,000 chest X-rays and 30,000 unique patients [42]. It originally contained 15 classes (14 diseases and 1 class for no finding). We split the dataset into two groups, where medical task 1 specifically tries to classify Atelectasis, Consolidation, Infiltration, Pneumothorax, and medical task 2 tries to classify Nodule, Mass, and Hernia. An example image taken from the dataset is shown in Figure 8 (a).
[…]
Figure 8: (a) Example figure from the NIH [42] dataset. (b) Pareto fronts found by brute-force direct search using 400 points and by MAP using 30 points. We randomly sampled 25 points from the predicted Pareto front by MAP. The resulting IGD is 0.016, and GD is 0.014.
F.2 Merging language model
We merged language models in French and Arabic, as well as Chinese and Japanese. Please refer to Table 7 and Figure 9. As we can see, the ground truth Pareto front is not in good shape. There are only a few points on the ground truth Pareto fronts which means few merged models could dominate the rest and the trade-off between the metrics of different languages might not be very significant. Even under this condition, our algorithm is still able to find out the Pareto fronts. We further try to merge ‘Arabic+French’ and ‘Chinese+Japanese’ in the nested scheme, Algorithm 2 with various preferences. However, the Pareto front usually only contains a single model which prevents us from predicting a Pareto front.
Table 7: Llama-3 fine-tuned models merging
GD IGD GD+IGD
Arabic+French $0.023_{0.010}$ $0.035_{0.018}$ $0.058_{0.028}$
Chinese+Japanese $0.014_{0.013}$ $0.028_{0.017}$ $0.041_{0.026}$
[…]
Figure 9: (a) Amortized Pareto front on merging Llama-3 fine-tuned on French with Llama-3 fine-tuned on Arabic; (b) Amortized Pareto front on merging Llama-3 fine-tuned on Chinese with Llama-3 fine-tuned on Japanese
F.3 Additional experiment results on ResNet
We performed additional experiments on ResNet18 [20] on merging two models finetuned on CIFAR10 [32] and Flowers102 [41] and show the Pareto front obtained in Figure 10. Unlike ViT models which perform zero-shot classification, ResNet requires additional fine-tuning of the classification head after model merging. We demonstrate that our method still applies to those models.
Figure 10: Pareto front obtained for two ResNet18 models on CIFAR10 and Flowers 102. We perform additional finetuning of the classification head after merging the model weights.
Appendix G More details about algorithms
G.1 Algorithm 1
In this section, we extend the discussion on line 7 of the algorithm and on Remark 1. In (5), the optimization problem is essentially minimizing the mean squared error (MSE) of a linear regression. The predictors are $\mathbf{C}=\text{concat}(c_{1}^{2},c_{2}^{2},\cdots,c_{N}^{2},c_{1}c_{2},c_{1}c_{3},\cdots,c_{T-1}c_{T},c_{1},c_{2},\cdots,c_{N},1)$, the variables are $(\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*})$, and the responses are $M_{n}(\bm{\theta}_{\text{m}}(\mathbf{c}))$.
We can certainly seek other designs to improve the accuracy and robustness of $(\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*})$. On one hand, we can use other loss functions like mean absolute error or Huber loss, while still using the linear regression. On the other, we can modify $\tilde{M}_{n}(\mathbf{c})$. If the metric $M_{n}$ has a specific bounded region, we could restrict the fitting interval. For example, if the metric is the accuracy or F1 score between 0 and 1, we could switch the definition to $\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv\text{sigmoid}(e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c})$. This is equivalent to a logistic regression with the same definition of predictors and responses as the linear regression. If the metric is loss, which ranges from 0 to positive infinity, then we could switch to $\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv\text{softplus}(e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c})$. In summary, one may extend (5) to
$\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*}=\arg\min_{\mathbf{A}_{n},\mathbf{b}_{n},e_{n}}\text{MSE}\left(M_{n}(\bm{\theta}_{\text{merge}}(\mathbf{c})),\sigma(\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n}))\right),$
and replace MSE and $\sigma=\mathbb{I}$ with other functions.
We emphasize that Algorithm 1 is an out-of-box plugin-like method. Many parts of it are kept generic: there are multiple ways to sample $\Omega$ in terms of the number of $\mathbf{c}$ and the style of sampling (Bayesian or not), thus creating a tradeoff between the quality of Pareto front and the computational time; the task vector can be computed on the more memory-capable CPU and possibly in a parameter-efficient manner (e.g. LoRA, Adapter, and BiTFiT): for example, computing $\mathbf{v}_{n}$ via the subtraction of two full 7B models requires $2\times 23=46$GB of memory capacity, but the memory cost for PEFT models may reduce to 16MB by only subtracting the non-frozen components; the for-loops in line 2 and 4 can be computed in parallel, possibly using a much smaller subset of the data, and thus enjoying a computational speed-up.
Remark 1 (quadratic surrogate model fitting)
In Algorithm 1, $\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*}$ in (5) can be readily solved by off-the-shelf methods, such as close-form solution (if applicable), LBFGS, Newton’s method and gradient descent, since this is a convex problem. In Appendix G, we generalize the ways to learn the coefficients in (4), besides minimizing the mean square error in (5).
G.2 Algorithm 2
In this section, we explain the operations of the algorithm in Figure 3 in details. Here task 1 to 8 is Cars, GTSRB, DTD, SUN397, Resisc45, and SVHN. If we minimize (3) without the nested merging, we would need to estimate $\mathbf{A}_{1},...,\mathbf{A}_{8}\in\mathbb{R}^{8\times 8}$, with hundreds of $\mathbf{c}$.
With the nested merging, for the first round, we merge $(\bm{\theta}_{ft}^{1},\bm{\theta}_{ft}^{2})$into $\bm{\theta}_{\text{merge}}^{1,2}$, thus approximating $\mathbf{A}_{1}$ and $\mathbf{A}_{2}\in\mathbb{R}^{8\times 8}$ by $\mathbf{A}_{1}[1:2,1:2]$ and $\mathbf{A}_{2}[1:2,1:2]\in\mathbb{R}^{2\times 2}$, respectively. That is, we only care about the interference between task 1 and 2, but not task 1 and 5. Similarly, we merge $(\bm{\theta}_{ft}^{3},\bm{\theta}_{ft}^{4})$into $\bm{\theta}_{\text{merge}}^{3,4}$, and $(\bm{\theta}_{ft}^{5},\bm{\theta}_{ft}^{6})$into $\bm{\theta}_{\text{merge}}^{5,6}$. Next, we merge $(\bm{\theta}_{\text{merge}}^{1,2},\bm{\theta}_{\text{merge}}^{3,4})$into $\bm{\theta}_{\text{merge}}^{1,2,3,4}$, and finally into $\bm{\theta}_{\text{merge}}^{1,2,3,4,5,6,7,8}$.
G.3 Algorithm 3
Algorithm 4 is a detailed version of Algorithm 3. Figure 11 includes illustration of our discretization method (how we create bins) in 2D and 3D decision variable ($\mathbf{c}$) space.
Figure 11: (a) Discretizing of two task scaling coefficients along the angular dimension in 2D polar coordinate system; (b) Discretizing of three task scaling coefficients along the angular dimensions in 3D spherical coordinate system;
Algorithm 4 Bayesian Adaptive of Surrogate Model
0:  Number of iterations $J$, Buffer $\mathcal{B}$, Pretrained model $\bm{\theta}_{\textup{pre}}$, Task vectors $\mathbf{v}_{n}$, Evaluators for task $N$, $M_{n}(\cdot)$, Discretization bin number $K$, sample size for every iteration $n_{j}$, $j=0$ to $J$, Bootstrap dropping rate $\alpha=20\%$, Bootstrap sampling number $Q=30$.
1:  $\mathcal{B}\leftarrow\emptyset$
2:  for $j=0$ to $J$ do
3:     if $j=0$ then
4:        Sample $n_{0}$ scaling coefficients $\{\mathbf{c}_{i}\}_{i=1}^{n_{j}}$ from $U([0,1]^{N})$
5:     else
6:        Sample $n_{j}$ scaling coefficients $\{\mathbf{c}_{i}\}_{i=1}^{n_{j}}$ based on the posterior distribution
7:     for $i=0$ to $n_{j}$ do
8:        Merge the model $\theta_{m}(\mathbf{c}_{i})=\bm{\theta}_{\textup{pre}}+\mathbf{c}_{i}\cdot\mathbf{v}_{n}$
9:        Evaluate $m_{n,i}=M_{n}(\theta_{m}(\mathbf{c}_{i}))$
10:        $\mathcal{B}\leftarrow\mathcal{B}\cup\{(\mathbf{c}_{i},m_{n,i})\}$
11:     Fit the quadratic approximation surrogate model $\tilde{M}_{n}$ by learning $\mathbf{A}_{n}^{*},\mathbf{b}_{n}^{*},e_{n}^{*}$ in (5).
12:     Discretize the scaling coefficients along the angular dimensions in hyper-spherical coordinates (see figure 11 as examples)
13:     for $k=0$ to $K$ do
14:        Calculate the mean of $L2$ loss between $\tilde{M}_{n}(\mathbf{c}_{i})$ and ${M}_{t}(\mathbf{c}_{i})$, where $\mathbf{c}_{i}$ are in bin $k$, denoted as $\text{mean}_{k}$Bootstrap to estimate the standard deviation of the losses.}
15:        for $q=0$ to $Q$ do
16:           Randomly (uniformly) drop $\alpha$ scaling coefficient in bin $k$
17:           Calculate the mean of $L2$ loss between $\tilde{M}_{n}(\mathbf{c}_{i})$ and ${M}_{t}(\mathbf{c}_{i})$ with the rest points and denoted with $l_{q}$
18:        Calculate the standard deviation of the $\{l_{q}\}_{q=0}^{Q}$ and denoted as $\text{std}_{k}$
19:        $\text{score}_{k}=\text{mean}_{k}+\frac{1}{2}\text{std}_{k}$
20:     Calculate probability distribution across the discretized bins by $\text{score}_{k}$ as the posterior sampling strategy in the next round
21:  return
Appendix H Potential QA
Q1: Why nested-merging MAP does not give a complete Pareto front?
In nested-merging MAP, we merge the models in pair. For example, there are $model_{i}$ for task i as individual fine-tuned models. $i=1,2,3,4$. We first merge model 1 and model 2 given the preference of the user between task 1 and task 2. We then get $model_{1,2}$. At the same time, we merge model 3 and model 4 given the preference of the user between task 3 and task 4 and get $model_{3,4}$. Finally, we merge the $model_{1,2}$ and $model_{3,4}$ given the preference of user to get $model_{1,2,3,4}$. We output the $model_{1,2,3,4}$ as the output merged model of the algorithm. Please note that the merging order in the algorithm should be decided by the loss function clustering. It is a heuristic decision and does not always dominate other merging orders. Thus, we choose not to emphasize the contribution of this order. The practitioner may use any order they think can be helpful on the merging.
Q2: It seems nested-merging MAP performs worse than MAP. What is the meaning of it?
In theory, the surrogate model can be easily fitted when the number of tasks is low. When it is high (e.g. 8), the fit of the surrogate model can no longer be a near-perfect approximation (please find the R2 in Table 2). Thus, even if the NMMAP can only find the suboptimal solution, it is still comparable to MAP Algorithm 1 according to Table 4. We understand in general that NMMAP does not find the global optimum, but please kindly keep in mind that even if their performance is comparable, NMMAP takes way less computation of evaluation.
Q3: Why nested-merging MAP does not output the optimal solution?
The solution found by nested-merging MAP (NMMAP) is indeed suboptimal given the limited search space compared with merging all models at once. However, it does not mean that it is not useful in all situations. When the number of tasks is high, it has comparable performance to MAP while consuming much fewer computations on evaluation.
Q4: The authors assume a setting where evaluation models are computationally expensive to query. But is this the reality?
[…]
Q5: Why does quadratic approximation have a lower cost?
For the MAP algorithm (Algorithm 1), the time complexity is the same as what we mentioned above, what is different is that we fitted an approximation of the evaluation metrics. We only need the scaling coefficient vectors to input to a quadratic function model to be able to get the estimated evaluation score. Run the quadratic function once take only $3.91\times 10^{-6}s\pm 894\times 10^{-9}s$. And thus, evaluating $2500$ times takes $<2500\times 4\times 10^{-6}s=10^{-2}s$.
Q6: Why not compare to Git-Rebasin [1]?
We didn’t compare our method to Git Re-Basin because their study focuses on the scenarios of merging the models from different initializations, whereas our background works on the same initialization of different fine-tuned models.
Q7: How do you deal with gradient and Hessian in the second-order Taylor expansion?
Notations:
  • $p$ as the number of parameters in the pre-trained model (also the number of parameters in each task vector).
  • $\mathbf{V}$ is the matrix of task vectors of different N tasks. Thus, $\mathbf{V}\in\mathbb{R}^{p\times N}$.
  • $\mathbf{c}$ is the scaling coefficient vector $\in\mathbb{R}^{N}$.
  • $M_{n}$ is the metric (e.g. accuracy) for the task n.
$M_{n}(\mathbf{c})$ $=\underbrace{M_{n}(\bm{\theta}_{\textup{pre}})}_{\in\mathbb{R}}+\underbrace{\nabla M_{n}(\bm{\theta}_{\textup{pre}})^{\top}}_{\in\mathbb{R}^{1\times p}}\underbrace{\mathbf{V}}_{\in\mathbb{R}^{p\times N}}\underbrace{\mathbf{c}}_{\in\mathbb{R}^{N\times 1}}+\frac{1}{2}\underbrace{(\mathbf{V}\mathbf{c})^{\top}}_{\in\mathbb{R}^{1\times p}}\underbrace{\mathbf{H}_{n}(\bm{\theta}_{\textup{pre}})}_{\in\mathbb{R}^{p\times p}}\underbrace{\mathbf{V}\mathbf{c}}_{\in\mathbb{R}^{p\times 1}}+\underbrace{R_{n}}_{\in\mathbb{R}}$ (6)
$=\underbrace{e_{n}}_{\in\mathbb{R}}+\underbrace{\mathbf{b}_{n}^{\top}}_{\in\mathbb{R}^{1\times N}}\underbrace{\mathbf{c}}_{\in\mathbb{R}^{N\times 1}}+\underbrace{\mathbf{c}^{\top}}_{\in\mathbb{R}^{1\times N}}\underbrace{\mathbf{A}_{n}}_{\in\mathbb{R}^{N\times N}}\underbrace{\mathbf{c}}_{\in\mathbb{R}^{N\times 1}}$ (7)
$\tilde{M}_{n}(\mathbf{c};\mathbf{A}_{n},\mathbf{b}_{n},e_{n})\equiv e_{n}+\mathbf{b}_{n}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{n}\mathbf{c}$
where
$\mathbf{A}_{n}=\mathbf{V}^{\top}\mathbf{H}_{n}(\bm{\theta}_{\textup{pre}})\mathbf{V}\in\mathbb{R}^{N\times N},\mathbf{b}_{n}=\mathbf{V}^{\top}\nabla M_{n}(\bm{\theta}_{\textup{pre}})\in\mathbb{R}^{N},e_{n}=M_{n}(\bm{\theta}_{\textup{pre}})+R_{n}$ (9)
Please notice that $\mathbf{A}$ is a symmetric matrix. Specifically, when the number of tasks is 2, we have:
$\tilde{M}_{1}(\mathbf{c};\mathbf{A}_{1},\mathbf{b}_{1},e_{1})$ $\equiv e_{1}+\mathbf{b}_{1}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{1}\mathbf{c}$ (10)
$=\frac{1}{2}A_{1,11}c_{1}^{2}+A_{1,12}c_{1}c_{2}+\frac{1}{2}A_{1,22}c_{2}^{2}+b_{1,1}c_{1}+b_{1,2}c_{2}+e_{1}$ (11)
$\tilde{M}_{2}(\mathbf{c};\mathbf{A}_{2},\mathbf{b}_{2},e_{2})$ $\equiv e_{2}+\mathbf{b}_{2}^{\top}\mathbf{c}+\frac{1}{2}\mathbf{c}^{\top}\mathbf{A}_{2}\mathbf{c}$ (12)
$=\frac{1}{2}A_{2,11}c_{1}^{2}+A_{2,12}c_{1}c_{2}+\frac{1}{2}A_{2,22}c_{2}^{2}+b_{2,1}c_{1}+b_{2,2}c_{2}+e_{2}$ (13)
We don’t calculate gradient or Hessian to get $\mathbf{A}_{1}$, $\mathbf{A}_{2}$, $\mathbf{b}_{1}$, $\mathbf{b}_{2}$, $e_{1}$ and $e_{2}$. We use linear regression to estimate them. How? Given a $(c_{1},c_{2})$ pair, we can define a merged model $\bm{\theta}_{\text{merge}}(c_{1},c_{2})$, we evaluate the merged model on task 1 and task 2 to get the metrics (e.g. accuracy). Given 20 pairs of $(c_{1},c_{2})$, we would be able to evaluate get 20 corresponding $(M_{1},M_{2})$ which are metric for task 1 and metric for task 2. Thus, we can fit the surrogate model $\tilde{M}_{1}(\mathbf{c};\mathbf{A}_{1},\mathbf{b}_{1},e_{1})$ and $\tilde{M}_{2}(\mathbf{c};\mathbf{A}_{2},\mathbf{b}_{2},e_{2})$.